SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Operation Meeting (01/22/08)

Attendees: Rainer Bartoldus, Su Dong, Mark Tibbetts, Rahmat Rahmat, Debbie Bard, Kim Hojeong, Steve Sekula.

Meeting Information

 Date      :  January 22nd, 2008
 Time      :  2:00pm - 3:30pm PST
 SLAC Room :  Madrone conf room, ROB 
 
 Dial-in   :  1-510-665-5437, pass code 6649#

--- Minutes ---

Operations

  • DCC board for TSF-X replaced by Selina and Hojeong on Friday around 5pm. The old board is in teststand.
  • This morning there was a power dip, and had many "spurious" errors on ROM log. Other than that, the ROM log hasn't shown much errors since. Maybe more stable but too soon to judge.
  • As Rainer suggested, I ask DIRC if they see any change in their bad timing after the board change. Since the problem was due to trickling, the effect is expected to be small. Jose will let us know if he sees any effect.

EMT

  • Emailed on hot tower. It was a short run with high background. No EMT problem.
  • During calibration after power dip, EMC calibration failed, and EMT calibration seems never tried. Mark will update the instruction, in case EMC calibration fails, perform a stand alone EMT calibration.

DCT

GLT

Online

L3

DQM

  • The ratios of Lumi Bhabha (Lumi GamGam, Lumi multihadron) over mu pair went down during 3S run. Is it expected? Look at p.121 of this file
  • The ratio of recorded Luminosity (from ir2runs) over MuPair counting (from BtaLumiCountnPlot) in Y(3S) run is lower. Look at HN posting
  • The change maybe due to acceptance change (smaller) and/or muon crosssection at 3S resonance (bigger). The acceptance change was introduced on Jan 5th. Looking before and after Jan 5th to compare would be useful.

Trickle injection monitoring

Trigger Configuration

  • Regarding Y(3S)->Y(1S)pi0pi0 triggering, look at the following posts:
    • Trigger Configuration Stdy for Y(3S) (link) by Su Dong
    • L3 effect (link) by Rainer
    • Recoil mass plots (link) by Steve

AOB


This page was created by Kim Hojeong; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.

Last Modified: Jan 22nd, 2008.